3.In how many ways can the letters in the word TEXAS be arranged if
a.) All the letters are used
b.) 2 are used at a time
c.) 4 are used at a time

Answer by ikleyn(52790)

.

(a)  5! = 5*4*3*2*1 = 120 arrangements (permutations).


(b)  5*4 = 20 2-letter words, if 2 letters are used at a time.


(c)  5*4*3*2 = 120 4-letter words, if 4 letters are used at a time.

Solved.

-------------

Since the order is important,  this problem is on  PERMUTATIONS.


On  Permutations,  see introductory lessons
    - Introduction to Permutations
    - PROOF of the formula on the number of Permutations
    - Simple and simplest problems on permutations
    - Special type permutations problems
in this site.

Also,  you have this free of charge online textbook in ALGEBRA-II in this site
    - ALGEBRA-II - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic  "Combinatorics: Combinations and permutations".


Save the link to this textbook together with its description

Free of charge online textbook in ALGEBRA-II
https://www.algebra.com/algebra/homework/complex/ALGEBRA-II-YOUR-ONLINE-TEXTBOOK.lesson

into your archive and use when it is needed.